# von Neumann entropy and quantum version of thermodynamic entropy

### v1

@@ Abstract
The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history. In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be either avoided or addressed convincingly. In a broader context, we analyze here whether and when the VN entropy is the same or equivalent to the TH entropy. For a thermalized isolated or open system, the VN entropy for an appropriately chosen density operator is the same as the quantum version of, respectively, the Boltzmann or Gibbs entropy (these latter entropies are equivalent to the TH entropy for large thermalized systems). Since the quantum thermalization is essentially defined for a subsystem of a much larger system, it is important to investigate if the VN entropy of a subsystem is equivalent to its TH entropy. We here show that the VN entropy of a small subsystem of a large thermalized system is proportional to the quantum statistical entropy (Boltzmann’s version) of the subsystem. For relevant numerical results, we take a one-dimensional spin-1/2 chain with next-nearest neighbor interactions.
@@ I Introduction
Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6]. For isolated system, Boltzmann gave formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates). For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms for probability distribution function in phase space. The quantum versions of these statistical entropies (Boltzmann’s and Gibbs’) are somewhat intuitively clear [7, 5]: for an isolated quantum system, the entropy is given in terms of the dimension of the relevant subspace of the system’s Hilbert space while for an open quantum system, the entropy can be calculated from the partition function (which depends on the eigenvalues of the appropriate Hamiltonian of the system). Question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system. We will address this issue separately in this work.
@@
From the thermodynamical considerations, von Neumann (VN) gave a formula for entropy of a quantum ensemble (representing a given mixed state) [8]. This entropy, known as VN entropy, attracted a lot of attention and criticism in regards to its connection to the TH entropy. We show in this work that, for isolated and open quantum system or for small part of a large quantum system, the VN entropy is equivalent to the TH entropy if the quantum system is large and in thermal equilibrium. We largely exclude from our discussions the quantum systems which do not thermalize or systems not in equilibrium.
@@
It is worth mentioning here the logical steps we follow in this work while establishing the said equivalence between the entropies and finding the conditions under which the equivalence works. First, identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: dimension of relevant subspace in place of volume of accessible phase-space, and in Gibbs’ case: energy eigenvalues in place of continuous classical energy). In the last step of understanding the equivalence, one checks if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e. the reinterpreted statistical entropy in terms of quantum variables). To better understand the equivalence, one should also check if the the VN entropy inherits the important properties of the equilibrium TH entropy. Accordingly, whereever appropriate, we briefly deliberate, if the VN entropy is a state function, additive and whether it increases for isolated system due to an irreversible process.
@@
We organize our paper in the following way. In the next section (Sec. II), we review the thermodynamic and statistical entropies, and subsequently, discuss the suitable quantum versions of these statistical entropies. In Sec. III, we discuss von Neumann entropy and some reservations about the entropy while it comes to representing the thermodynamic entropy. Next, in Sec. IV, we check for three different cases (isolated system, open system and subsystem) if the von Neumann entropy is equivalent to the thermodynamic entropy. We conclude our work in Sec. V.
@@ II.1 Entropy: from Clausius to Boltzmann and Gibbs
The Clausius theorem or the Clausius inequality states that [4], for any cyclic transformation or thermodynamic process, $\oint\frac{d\bar{}\hskip 0.70004ptq}{T}\leq 0$, where $d\bar{}\hskip 1.00006ptq$ is the heat increment received by the system at temperature $T$. Here the equality sign works only for a reversible cycle and this helps us to define a thermodynamic (TH) state function, called entropy. Streaky speaking, this state function (thermodynamic entropy, $S_{TH}$) is well defined only for a thermodynamic system in equilibrium. If a system in equilibrium at temperature $T$ receives $d\bar{}\hskip 1.00006ptq$ amount of heat reversibly and subsequently reaches an equilibrium again, then the difference in entropy between these two equilibrium phases would be $dS_{TH}=\frac{d\bar{}\hskip 0.70004ptq}{T}$.
@@
II.2 Conditions of equivalence between thermodynamic and statistical entropy
@@
To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence. A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [6]. Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat. Although the systems exchange heat, total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems. In the physical setup to establish the equivalence, the main premise is that a thermodynamic system in equilibrium with a given energy can be in any of a certain number of microstates with an equal probability (the principle of equal a priori probabilities). Here the number of accessible microstates is a function of the system’s energy. If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then at this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$. Here we assume that the two systems can independently be in any of their accessible microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored). The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values. As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates. When the systems are thermodynamically large, $M(E)$ will be almost equal to $\Omega(\overline{E}_{1})$, i.e., $M(E)\approx\Omega(\overline{E}_{1})$, where $E^{\prime}_{1}=\overline{E}_{1}$ maximizes the function $\Omega(E^{\prime}_{1})$. This means that the composite system, after settling in a new equilibrium state, will spend most of its time in the microstates corresponding to the energy distribution $\overline{E}_{1}$ and $E-\overline{E}_{1}$ for the first and second systems, respectively.
@@
We stress here that, when the system is in non-equilibrium, it would not be equally probable to be in different microstates (at a given time, a subset of microstates may be more probable than the other subsets). Only in equilibrium, as per the premise, the system is equally probable to be in any of the accessible microstates. Now since in equilibrium the composite system is equally probable to be any of the $M(E)$ microstates, which is approximated by $\Omega(\overline{E}_{1})$, a condition for the new equilibrium between the systems can be established by a maximization condition of $\Omega(E^{\prime}_{1})$ (or, in practice, $\ln\Omega(E^{\prime}_{1})$). After comparing this condition with a standard thermodynamic relation (namely, $\frac{1}{T}=\left(\frac{\partial S}{\partial E}\right)_{N,V}$), one infers that the thermodynamic entropy of the (first) system is equal to the BN entropy, i.e., $k_{B}\ln\Omega_{1}(\overline{E}_{1})$.
@@
From the above discussion, we conclude that the TH entropy and the BN entropy of an isolated system is equivalent ($S_{TH}\equiv S_{BN}$) if the system is large (so that the temperature $T$ can be defined and $M(E)$ can be approximated by $\Omega(\overline{E}_{1})$) and the system is in thermal equilibrium (so that the principle of equal a priori probabilities becomes applicable).
@@ II.3 Quantum version of statistical entropy
It is not straightforward how one can reinterpret the BN or GB entropy in terms of quantum variable(s). This is because, one of the conditions for equivalence, namely the principle of equal a priori probabilities, needs re-examination. To understand this, we recall that, a classical system moves from one microstate to another due to its time evaluation (each point in a system’s equal-energy trajectory is a microstate and they are statistically equally probable). On the other hand, what is a microstate for a quantum system? It is tempting to think of the eigenstates of the system’s Hamiltonian as the microstates, as one can then refer to all the eigenstates withing a narrow energy window to introduce the principle of equal probabilities for a quantum system. The main problem with this assumption is that the system does not move from one eigenstate to another due to its time evaluation. In the following we discuss how one may resolve this problem.
@@
If we consider an isolated large quantum system with size $N$ (the ‘size’ can be interpreted as the number of particles or constituents of the system), the average gap between any two consecutive eigenvalues will be exponentially small with the system size. We note that the (average) density of states is proportional to the Hilbert space dimension ($D$) and the average gap ($\delta$) is inversely proportional to the density of states. Since $D\propto e^{cN}$, we have $\delta\propto e^{-cN}$, where $c$ is a positive constant. Thus, if $N$ is large enough, it will be practically impossible to prepare the quantum system in a pure state. In other words, because of our said inability, we only can study a system which can be in any of the many closely spaced eigenstates with some probability distribution. Here we may apply the principle of equal a priori probabilities and assume the probability distribution to be uniform in the equilibrium (because, we do not have any specific reason to believe that one state will be preferred over the others in the same energy range).
@@
There are other ways one can justify this equal probabilities principle for a quantum system. For example, due to some (weak) interaction in past, currently the isolated system can be in an entangled state with its environment. This effectively puts the isolated system in a mixed state, i.e., the system can be in any of many microstates (eigenstates) corresponding to a particular energy scale. The associated probability distribution, as per the principle, can be taken to be uniform. Yet another way to justify the principle is the following. Let $H_{0}$ be the main Hamiltonian of the noninteracting system and $H^{\prime}$ represents the effective (weak) interaction among the particles (taken to be perturbation to the main Hamiltonian). Hence the total Hamiltonian describing the system is $H=H_{0}+H^{\prime}$. Now if we take eigenstates of $H_{0}$ to be the microstates, the system will make transitions from one to other microstates during time evaluation due the effect of the interaction $H^{\prime}$. If the effective interaction is taken to be weak, all the eigenstates of $H_{0}$ from a narrow energy window can be assumed to be equally probable, enabling us to use the results of statistical mechanics.
@@
Once we have agreement on quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies. To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets. As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
@@
To get the quantum version of Gibbs (qGB) entropy, in place of continuous energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function. Here we assume that, the system’s interaction with the reservoir is weak (it had interaction with the reservoir in the past) so that we still can effectively describe the dynamics of the system by its Hamiltonian. As before, let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian. In the spirit of Gibbs distribution for classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. Subsequently, we can write the quantum version of the GB entropy in the following way:
@@
This formula attracted both the applause and criticism alike [10, 11, 12, 13, 14]. In the following two subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies. Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or are unfounded.

### v2

@@ Abstract
The debate whether the von Neumann (VN) entropy is a suitable quantum version of the thermodynamic (TH) entropy has a long history. In this regard, we briefly review some of the main reservations about the VN entropy and explain that the objections about its time-invariance and subadditivity properties can be addressed convincingly. In a broader context, we here analyze in detail whether and when the VN entropy is the same or equivalent to the TH entropy. We find that the VN entropy of a large isolated system is equivalent to its TH entropy if the system is in a mixed state described by a microcanonical (MC) ensemble. This equivalence also holds for open system and subsystem if the full system (i.e. system-bath or subsystem-bath combination) is in a mixed state described by an MC ensemble or in a typical pure state (or if the Eigenstate Thermalization Hypothesis, ETH, holds for the full system). Unlike isolated and open system, demonstrating the equivalence for a subsystem is not straightforward. We take a one-dimensional spin-1/2 model to numerically show that, if the full system is in a pure state, then the VN entropy of a subsystem is equivalent to its TH entropy if the ETH holds for the full system. We also demonstrate that this equivalence always works (irrespective of whether the ETH holds) if the full system is in a mixed state describe by an MC ensemble.
@@ I Introduction
Boltzmann, and later Gibbs, showed us how the TH entropy can be calculated from the microscopic details using statistical approaches [4, 6]. For an isolated system, Boltzmann gave a formula (actually Planck first wrote down the explicit expression for the Boltzmann entropy) for the TH entropy in terms of the volume of the relevant phase space (or the number of associated microstates). For the open system, Gibbs on the other hand, using his ensemble theory, provided the formula for the TH entropy in terms of an appropriate probability distribution function in phase space. Although the quantum versions of these statistical entropies for the isolated and open systems in equilibrium are somewhat intuitively clear [7, 5], one needs convincing justifications, which we discuss in Sec. II. The question, however, remains on the quantum version of the TH entropy for a part (subsystem) of a quantum system. In this work, we will discuss this issue in detail.
@@
As noted, for a quantum system in thermal equilibrium, writing down the quantum version of the Gibbs (GB) entropy is somewhat easy (see, Secs. II.3 and III.1). But when a quantum system is in a generic mixed state and not in equilibrium, what would be the value of entropy? From thermodynamical considerations, von Neumann (VN) gave a formula for the entropy of a quantum ensemble (representing a given mixed state) [8]. This entropy, known as VN entropy, attracted a great deal of attention and a lot of criticism in regards to its connection to the TH entropy. We show in this work that, for isolated and open quantum system or for a subsystem of a (much) larger quantum system, the VN entropy is equivalent to the TH entropy if the quantum system (or subsystem) is large and in thermal equilibrium. We largely exclude from our discussions the quantum systems that do not thermalize or systems that are not in equilibrium. We here, though, deliberate on two main criticisms on the VN entropy, related to the time-invariance and subadditivity, and argue that these criticisms can be avoided or addressed convincingly.
@@
We outline here the logical steps we follow in this work while establishing the said equivalence between the VN and TH entropies and finding the conditions under which the equivalence works. First, we identify under what conditions the TH entropy is equivalent to the statistical entropy (either Boltzmann’s or Gibbs’ version), then reinterpret the statistical entropy in terms of the quantum variables (in Boltzmann’s case: the dimension of relevant subspace in place of the volume of accessible phase-space, and in Gibbs’ case: the energy eigenvalues in place of the continuous classical energy). In the last step of establishing the equivalence, we check if the VN entropy of an appropriate density matrix is equivalent to the quantum version of the statistical entropy (i.e. the reinterpreted statistical entropy in terms of quantum variables). To better understand the equivalence, one should also check if the VN entropy inherits the important properties of the equilibrium TH entropy. Accordingly, wherever appropriate, we briefly deliberate if the VN entropy is a state function, has the additivity property and whether it increases for isolated system due to an irreversible process.
@@
We organize our paper in the following way. In the next section (Sec. II), we review the thermodynamic and statistical entropies, and discuss the condition of equivalence between these entropies. This is followed by a discussion on the suitable quantum versions of the statistical entropies. In Sec. III, we discuss von Neumann entropy and some reservations about the entropy when it comes to representing the thermodynamic entropy. Next, in Sec. IV, we check for three different cases (isolated system, open system and subsystem) if the von Neumann entropy is equivalent to the thermodynamic entropy. We conclude our work in Sec. V.
@@ II.1 Entropy: from Clausius to Boltzmann and Gibbs
The Clausius theorem or the Clausius inequality states that [4], for any cyclic transformation or thermodynamic process, $\oint\frac{d\bar{}\hskip 0.70004ptq}{T}\leq 0$, where $d\bar{}\hskip 1.00006ptq$ is the heat increment received by the system at temperature $T$. Here the equality sign works only for a reversible cycle and this helps us to define a thermodynamic (TH) state function, called entropy. Strictly speaking, this state function (thermodynamic entropy, $S_{TH}$) is well defined only for a thermodynamic system in equilibrium. If a system in equilibrium at temperature $T$ receives $d\bar{}\hskip 1.00006ptq$ amount of heat reversibly and subsequently reaches an equilibrium again, then the difference in entropy between these two equilibrium phases would be $dS_{TH}=\frac{d\bar{}\hskip 0.70004ptq}{T}$. In contrast, if the process is irreversible, then $dS_{TH}>\frac{d\bar{}\hskip 0.70004ptq}{T}$. This shows, for any isolated system undergoing a process ($d\bar{}\hskip 1.00006ptq=0$), the change in its entropy is $dS_{TH}\geq 0$.
@@
II.2 Conditions of equivalence between thermodynamic and statistical entropies
@@
To understand the conditions under which the equivalence works, we outline here a method to establish such equivalence. A common way of showing the equivalence between the TH and BN entropies of a system with a given energy is to bring the system with energy $E_{1}$ in contact with another system with energy $E_{2}$ [6]. Before the contact, the systems were in thermal equilibrium and after the contact they will again come to a common equilibrium via exchange of heat. Although the systems exchange heat, the total energy ($E=E_{1}+E_{2}$) of the composite system remains constant; here one ignores the interaction energy between the systems. In the physical setup to establish the equivalence, one generally employs the principle of equal a priori probabilities. According to this premise, a thermodynamic system in equilibrium with a given energy can be found in any of the available microstates with equal probability. Here the number of available microstates is a function of the system’s energy. If $\Omega_{1}(E^{\prime}_{1})$ ($\Omega_{2}(E^{\prime}_{2})$) is the number of microstates of first (second) system with energy $E^{\prime}_{1}$ ($E^{\prime}_{2}$), then in this energy distribution, the number of microstates of the composite system would be $\Omega(E^{\prime}_{1})=\Omega_{1}(E^{\prime}_{1})\Omega_{2}(E-E^{\prime}_{1})$. We here assume that the two systems can independently be in any of their available microstates (this assumption is related to our assumption that the interaction energy between the two systems can be ignored). The total number of microstates of the composite system with energy $E$ is $M(E)=\sum_{E^{\prime}_{1}}\Omega(E^{\prime}_{1})$, where the sum is over all possible $E^{\prime}_{1}$ values (i.e., over all possible energy distributions). As per the premise, the composite system in equilibrium is equally likely to be in any of the $M(E)$ number of microstates. When the systems are thermodynamically large, $M(E)$ will be almost equal to $\Omega(\overline{E}_{1})$, i.e., $M(E)\approx\Omega(\overline{E}_{1})$, where $E^{\prime}_{1}=\overline{E}_{1}$ maximizes the function $\Omega(E^{\prime}_{1})$. This means that the composite system, after settling in a new equilibrium state, will spend most of its time in the microstates corresponding to the energy distribution $\overline{E}_{1}$ and $E-\overline{E}_{1}$ for the first and second systems, respectively.
@@
We stress here that, when the system is in non-equilibrium, it would not be equally probable to be in different microstates (at a given time, a subset of microstates may be more probable than the other subsets). Only in equilibrium, as per the premise, the system is equally probable to be in any of the accessible microstates. Now since in equilibrium, the composite system is equally probable to be found in any of the $M(E)$ microstates, which is approximated by $\Omega(\overline{E}_{1})$, a condition for the new equilibrium between the systems can be established by a maximization condition of $\Omega(E^{\prime}_{1})$ (or, in practice, $\ln\Omega(E^{\prime}_{1})$). After comparing this condition with a standard thermodynamic relation (namely, $\frac{1}{T}=\left(\frac{\partial S_{TH}}{\partial E}\right)_{N,V}$), one infers that the thermodynamic entropy of the (first) system is equal to the BN entropy, i.e., $k_{B}\ln\Omega_{1}(\overline{E}_{1})$.
@@
From the above discussion, we conclude that the TH entropy and the BN entropy of an isolated system are equivalent ($S_{TH}\equiv S_{BN}$) if the system is large (so that the temperature $T$ can be defined and $M(E)$ can be approximated by $\Omega(\overline{E}_{1})$) and the system is in thermal equilibrium (so that the principle of equal a priori probabilities becomes applicable).
@@ II.3 Quantum version of statistical entropy
Intuitively it is not very difficult to guess the quantum version of a statistical entropy. However, it is not straightforward to see how the conditions discussed in the preceding subsection are satisfied for a quantum system. For example, one of the conditions for equivalence, namely the principle of equal a priori probabilities, needs re-examination. To understand this, we recall that, a classical system moves from one microstate to another due to its time evaluation (each point in a system’s equal-energy trajectory is a microstate and they are statistically equally probable). On the other hand, what is a microstate for a quantum system? It is tempting to think of the eigenstates of the system’s Hamiltonian as the microstates, as one can then refer to all the eigenstates withing a narrow energy window to introduce the principle of equal probabilities for a quantum system. The main problem with this assumption is that the system does not move from one eigenstate to another due to its time evaluation. In the following we discuss how one may resolve this problem.
@@
If we consider an isolated large quantum system with size $N$ (the ‘size’ can be interpreted as the number of particles in or constituents of the system), the average gap between any two consecutive eigenvalues will be exponentially small with the system size. We note that the (average) density of states is proportional to the Hilbert space dimension ($D$) and the average gap ($\delta$) is inversely proportional to the density of states. Since $D\propto e^{cN}$, we have $\delta\propto e^{-cN}$, where $c$ is a positive constant. Thus, if $N$ is large enough, it will be practically impossible to prepare the quantum system in a pure state. In other words, because of our said inability, we only can study a system which can be in any of the many closely spaced eigenstates with some probability distribution. Here we may apply the principle of equal a priori probabilities and assume the probability distribution to be uniform in the equilibrium (because, we do not have any specific reason to believe that one state will be preferred over the others in the same energy range).
@@
There are other ways one can justify this equal probabilities principle for a quantum system. For example, due to some (weak) interaction in past, currently the isolated system can be in an entangled state with its environment. This effectively puts the isolated system in a mixed state, i.e., the system can be in any of many microstates (eigenstates) corresponding to a particular energy scale. The associated probability distribution, as per the principle, can be taken to be uniform. Yet another way to justify the principle is the following. Let $H_{0}$ be the main Hamiltonian of the noninteracting system and $H^{\prime}$ represents the effective (weak) interaction among the particles (taken to be perturbation to the main Hamiltonian). Hence the total Hamiltonian describing the system is $H=H_{0}+H^{\prime}$. Now if we take the eigenstates of $H_{0}$ to be the microstates, then the system will make transitions from one to other microstates during time evaluation due to the effect of the interaction $H^{\prime}$. If the effective interaction is taken to be weak, all the eigenstates of $H_{0}$ from a narrow energy window can be assumed to be equally probable, allowing us to use the results of statistical mechanics.
@@
Once we have agreement on the quantum version of the principle of equal a priori probabilities, it is not difficult to come up with quantum versions of the statistical entropies. To get the quantum version of the BN (qBN) entropy, we prepare a system in such a way that its energy lies in the narrow energy shell $(E,E+\Delta E]$, where $\Delta E$ is arbitrary but macroscopically small and microscopically large [5]. Let $\{E_{n}\}$ be the set of eigenvalues of the system Hamiltonian and {|$\{\ket{n}\}$ be the corresponding set of eigenkets. As per the premise, the system will be in any state |$\ket{n}$ with equal probability if $E_{n}\in(E,E+\Delta E]$. Subsequently, guided by our understanding of the BN entropy, we can write its quantum version ($S_{qBN}$) in the following way:
@@
To get the quantum version of the Gibbs (qGB) entropy, in place of continuous classical energy, we use energy eigenvalues of the system’s Hamiltonian to calculate the partition function. Here we assume that the system’s interaction with the reservoir is weak (or it had interaction with the reservoir in the past), so that we still can effectively describe the dynamics of the system by its Hamiltonian. As before, let $\{E_{n}\}$ be the set of eigenvalues of the Hamiltonian of the (open) system. In the spirit of the Gibbs distribution for the classical system, we can define the probability to find the system with energy $E_{n}$ as $p_{n}=\frac{1}{Q}e^{-\beta E_{n}}$, where $Q$ is the partition function, $Q=\sum_{n}e^{-\beta E_{n}}$. Here $\beta$ denotes the inverse temperature. Subsequently, we can write the quantum version of the GB entropy in the following way:
@@
We here briefly note that the probabilities $p_{n}$ are the eigenvalues of the density matrix $\rho_{\beta}=\frac{e^{-\beta H}}{\text{Tr}(e^{-\beta H})}$, where $H$ is the Hamiltonian of the (open) system. Using the density matrix $\rho_{\beta}$, the quantum Gibbs (qGB) entropy in Eq.2 can also be expressed in a basis-independent form, as discussed in Sec.III.1. It may be emphasized that the relevant density matrix is found here indirectly: starting from a classical (open) system, we first find the probability distribution and then reinterpreting the quantities in terms of the quantum variables, we get the density matrix. Instead, it is also possible to get the density matrix $\rho_{\beta}$ directly, starting with a quantum system [10, 11]. In this work, we prefer the first method since it directly reveals when and how the TH entropy is equivalent to the qGB entropy.
@@
This formula attracted both the applause and criticism alike [12, 13, 14, 15, 16]. In the following subsections, we first discuss two different perspectives to understand how one can arrive at the formula and what it actually quantifies. Next we highlight some of the major criticisms on the applicability of the formula and explain our views that many of such criticisms can be avoided or can be addressed convincingly.